OpenStudy (anonymous):

Q 1 Write out the expansion of (1 - 2x)7 . Q 2 Are the lines and parallel, perpendicular, or neither? Q 3 Find the exact value of Cos 300o .

OpenStudy (anonymous):

7-14x cos 300=1/2

OpenStudy (anonymous):

Are the line y=2/3x+7and 3y=2x-9 parallel, perpendicular, or neither? i canot write properly before because it cannt write in message so plz solve this question.

OpenStudy (anonymous):

r u online mam..?

OpenStudy (anonymous):

parallel